The Caraway Seed Company grows heirloom tomatoes and sells their seeds. The heirloom tomato plants are preferred by many growers for their superior flavor. At the end of the most recent year the firm had current assets of $48,900​, net fixed assets of $249,000​, current liabilities of $29,600​, and​ long-term debt of $100,500.

a.  Calculate​ Caraway's stockholders' equity.

b. What is the​ firm's net working​ capital?

c.  If​ Caraway's current liabilities consist of  $21,200 in accounts payable and $8,400 in​ short-term debt​ (notes payable), what is the​ firm's net working​ capital?
Answer #1

a. Caraway's Stock Holders's equity =Net Fixed assets +Current assets -Current liabilities -Long term debt
=249000+48900-29600-100500 =167800

b. Firm's net working capital =Current Assets -Current Liabilities =48900-29600 =19300

c. Net Working capital will remain same at 19300.